OpenStudy (anonymous):

Fan and Medal? Can someone explain how to solve What is the equation in point-slope form of the line passing through (3, 6) and (−2, 1)? y + 6 = 1(x + 3) y + 1 = −1(x + 2) y − 6 = −1(x + 3) y − 6 = 1(x − 3)

OpenStudy (anonymous):

do you know the general equation for point slope form?

OpenStudy (anonymous):

Afraid not. :s

OpenStudy (anonymous):

\[y-y_{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Is it 6 - 1 / 3 - (-2)

OpenStudy (anonymous):

that will give you the slope (m)

OpenStudy (anonymous):

so then I use y = mx + b

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

y=mx+b is standard form. the question wants point-slope form

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

\[(x _{1},y _{1}) = (3,6) or (-2,1)\]

OpenStudy (anonymous):

what is the slope (m)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):